A 8.6 percent coupon (paid semiannually) bond, with a $1,000 face value and 10 years remaining to maturity. The bond is selling at $915.

value: 25.00 points Calculate the yield to maturity on the following bonds. a. A 8.6 percent coupon (paid semiannually) bond, with a $1,000 face value and 10 years remaining to maturity. The bond is selling at $915. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161)) Yield to maturity % per year b. An 5.7 percent coupon (paid quarterly) bond, with a $1,000 face value and 10 years remaining to maturity. The bond is selling at $911. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161)) Yield to maturity % per year c. An 7.7 percent coupon (paid annually) bond, with a $1,000 face value and 8 years remaining to maturity. The bond is selling at $1,061. (Do not round intermediate calculations. Round your answer to 3 decimal places. (e.g., 32.161)) Yield to maturity % per year
Answer #1

a)

Number of periods = 10 * 2 = 20

Coupon = (0.086 * 1,000) / 2 = $43

Yield to maturity = 9.963%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1000, PV -915, N 20, PMT 43, CPT I/Y

b)

Number of periods = 10 * 4 = 40

Coupon = (0.057 * 1,000) / 4 = $14.25

Yield to maturity = 6.942%

Keys to use in a financial calculator: 2nd I/Y 4, FV 1000, PV -911, N 40, PMT 14.25, CPT I/Y

c)

Number of periods = 8

Coupon = 0.077 * 1,000 = $77

Yield to maturity = 6.691%

Keys to use in a financial calculator: FV 1000, PV -1061, N 8, PMT 77, CPT I/Y